Fermi National Laboratory
 

Accelerator Update for June 23, 2006 - June 26, 2006

 

Friday June 23

The day shift began with Operations monitoring store 4794 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI) and Recycler operating normally, and with MiniBooNE and NuMI taking beam.

At 9:11 AM, a Booster vacuum burst closed valves, tripped off RF, and tripped off Linac. Operators had everything reset within twenty minutes.

At 10:25 AM, Operations reported that the first batch of the Booster beam was missing. It seemed to be a timing issue; investigations continued.

Tevatron (TeV) experts began a study around 11:13 AM. CDF and D0 turned off their detectors for this study.

At 11:40 AM, Operations reported beam position problems in the AP1 transfer line. Operations halted Pbar beam to investigate a possible ramp problem.

The TeV study ended at 1:21 PM.

Operators found a problem with a tuning program that was resolved with a reboot. Pbar resumed stacking at 1:42 PM.

Operations began preparing the Recycler at 5:25 PM to stash antiprotons. The stash ended and Pbar resumed stacking at 6:19 PM.

Operations terminated store 4794 at 7:10 PM.

Operations put the TeV into shot setup at 9:27 PM.

At 10:09 PM, the TeV suffered a sector F2 and F3 quench. Operations determined that beam was injected on top of other beam.

Saturday June 24

The midnight shift began with Operations monitoring the stash, with the TeV back in shot setup, with MI, Recycler, SY, and Booster operating normally, and with MiniBooNE and NuMI taking beam.

The Recycler began preparing to stash at 2:12 AM.

Operations established store 4797 at 2:14 AM with and initial luminosity of 74.1E30. Operators reported a collimator problem. They called an expert who had them pull the collimator out, recalibrate, and then rescrape.

The stash ended and Pbar resumed stacking at 2:54 AM.

At 8:02 AM, Operations reported ground water in the lower Linac. Operators notified the Duty Mechanic.

The Recycler began preparing to stash at 9:23 AM. The stash ended and Pbar resumed stacking at 10:06 AM.

At 10:41 AM, Operations reported stacking trouble. Operators thought it was the tuning program again, but after rebooting the program several times they called in an expert.

Operations established beam to the Meson train at 11:13 AM.

Operations established beam to the Meson MTest experiment around 12:40 PM.

At 5:05 PM, Operations discovered a Booster notcher problem. This might also be affecting Pbar. Operations called in a Booster expert.

MTest experimenters said that they were done with the beam at 6:07 PM.

The Recycler began preparing to stash at 7:53 PM. The stash ended around 8:50 PM.

At 8:56 PM, Operations reported that the MI RF tripped off due to a LCW pressure problem. An operator called an expert.

Pbar resumed stacking at 8:59 PM.

Sunday June 25

The midnight shift began with Operations monitoring store 4797 and the stash, with Pbar stacking, with normal MI, Recycler, SY, and Booster operations, and with MiniBooNE and NuMI taking beam.

At 3:59 AM, operators reported that they had problems turning on Ecool in preparation for the next shot. An operator called in an expert who managed to turn Ecool on by 4:57 AM.

Operation terminated store 4797 at 5:06 AM.

Operations put the TeV into shot setup at 5:18 AM.

Operations established store 4798 at 7:07 AM with an initial luminosity of 86.46E30.

The Recycler began preparing to stash at 8:10 AM.

At 8:33 AM, Operations reported that a Booster kicker (MKS08) had tripped off. Operators investigated and then called in an expert.

The stash ended at 8:35 AM.

Booster experts reported that they had repaired MKS08. Operations resumed sending beam to MTest, NuMI, MiniBooNE, and to Pbar at 11:56 AM.

The Recycler began preparing to stash at 5:30 PM. The stash ended and Pbar resumed stacking at 6:08 PM.

At 9:22 PM, the Tevatron Electron Lens tripped off. It reset okay.

Monday June 26

The midnight shift began with Operations monitoring store 4798 and the stash, with Pbar stacking, with normal MI, Recycler, SY, and Booster operations, and with MiniBooNE and NuMI taking beam.

The Recycler began preparing to stash at 12:44 AM. The stash ended and Pbar resumed stacking at 1:30 AM.

At 3:28 AM, Operations reported that the H- Source had tripped off. Operations called in experts. This problem held off all beam.

Beam returned at 8:04 AM.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/26/2006    email Fermilab